A pilot is required to fly directly from London, UK, to Rome, Italy.In order toarrive in Rome on time, the velocity of the plane relative to the ground must be 400 km/h [S43°E].If a wind is blowing with a velocity of 75 km/h [S], determine what the velocity of the plane relative to the air must be in order for the pilot to arrive in Rome on time.


1
Expert's answer
2022-03-10T10:18:20-0500

"u=v+v'\sin 43°=451~\frac{km}h."